You are considering investing RM65000 in new equipment. You estimate that the net cash flows will be RM12000 during the first year, but will increase by RM2500 per year the next year and each year thereafter. The equipment is estimated to have a 7-year service life and a net salvage value of RM5500 at that time. Assume MARR of 9%. a Calculate the annual capital cost CR (ownership cost) for the equipment. b.Determine the equivalent annual savings. c Is this a wise investment? YN Format: 74738 Format: 29879 Format A
Answer #1

PV of Cost

Particulars Amount Discounting factor PV
Investment 65000 1 65000
Net salvage value -5500 0.54703424 -3008.688347
Present value 61991.31165

Annual capital cost = 61991.31165/Annuity fator for 7 years of 9%= 61991.31165/5.032952835 = $12317.085752 i.e. $12317.09

PV of Cost savings

Year Cost savings Discounting factor PV of Cost savings
1 12000 0.91743119 11009.17431
2 14500 0.84167999 12204.3599
3 17000 0.77218348 13127.11916
4 19500 0.70842521 13814.29162
5 22000 0.64993139 14298.4905
6 24500 0.59626733 14608.54951
7 27000 0.54703424 14769.92461
Total 93831.90961

Equivalent annual savings = PV of savings/Annuity factor = 93831.90961/5.032952835 = $18643.5106157 i.e. $18643.51

NPV of the Investment = 93831.90961 - 61991.31165 = $31840.59796 i.e $31840.60

Since NPV is positive this is a wise Investment = Y